Mercury orbits the sun at a distance of 0.4 AU. What is mercury’s orbital period in earth

years? Keplers’s Third Law : T^2=d^3. T: orbital period in earth year. D: distance to sun

in astronomical units. 1 AU=150,000,000 or 1.5x10^8 km.


1
Expert's answer
2021-10-21T10:53:10-0400

Application of Kepler's third law gives us the following:

"T^2=d^3→
T=\sqrt{ d^3}=0.25\text{ years}."